OpenStudy (anonymous):

multiple choice. 72 is 75% of what number? A: 0.96 B:54 C:96 D:5400

OpenStudy (anonymous):

72 = 0.75 x solve

OpenStudy (anonymous):

96

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

I mean you could just divide 72 by 75% where 75% = 0.75 So 72/0,75 = 96 Or you could use a proportion in where you set up 75/100 = 72/x then you cross product. So now 75x = 7200 And then divide 75 from both sides to isolate x. So your answer, is again, 96. No matter what way you do it. Sorry the drawing isn't working right now so tell me if you need further explanation. :) x